Bob is 24 years younger than his father. In 2 years he will be half as old as his father. How old is each now?
Answer by stanbon(75887)

Bob is 24 years younger than his father. In 2 years he will be half as old as his father. How old is each now?
---------------
Let father's age be "f":
Bob's age is "f-24":
--------------------------
Equation:
(f-24+2) = (1/2)(f+2)
2(f-22) = f+2
2f-44 = f+2
f = 46 (father's age NOW)
f-24 = 22 (Bob's age NOW)
===========================
Cheers,
Stan H.